OpenStudy (anonymous):

covert to rectangular form r=sec@/(1+tan@)

OpenStudy (blockcolder):

Simplify the RHS first: \[\Large r=\frac{\frac{1}{\cos\theta}}{1+\frac{\sin\theta}{\cos\theta}}=\frac{1}{\cos\theta+\sin\theta}\\ \Large r(\cos\theta+\sin\theta)=1\\ \Large r\cos\theta+r\sin\theta=1\\ \Large x+y=1\]